# Control-Oriented Deep Space Communications For Unmanned Space Exploration

### v2

@@ I-A Background and Motivation
Communication plays a vital role in deep space exploration. The link between the spacecraft and the earth command center is the lifeline of the spacecraft. This connection provides tracking, telemetry and command (TT&C) services, which ensure the correct operation. According to the definition of the International Telecommunication Union , if the spacecraft-Earth distance is greater than $2\times 10^{6}$ km, the communication between the spacecraft and the earth belongs to deep space communication (DSC) [1]. Due to the great distance, DSC exhibits unique characteristics. The first is the large latency. Taking Mars as an example, the propagation delay from the perigee of Mars to Earth is 4 minutes, while from the apogee, it is 24 minutes [2]. The second is the low data rate. As a common sense, the signal power decays at least quadratically with distance. To ensure the same strength of the received signals, the transmit power of the spacecraft from Mars should be $9\times 10^{16}$ times stronger than that of a smart phone located 1 km away from the receiver [2]. However, constrained by size and weight limitations, the on-board base station (BS) carried by the spacecraft is less powerful than ground-based facilities. It can only transmit signals at very low power. Ground receivers must be equipped with large aperture antennas to capture the weak signals from space. Even so, the data rate between the spacecraft and the earth command center remains quite low, with only several megabits per second in NASA’s Deep Space Network (DSN) [3].
@@
With the growing curiosity on outer space, current space activities touch the fundamental problems of DSC. The sheer number of space missions is one factor. According to the United Nations, the number of space robots has grown exponentially in the last decade [4]. With advanced sensing techniques, new spacecrafts need a higher data rate to send high-resolution data back, which dramatically increases the burden of the DSN. Moreover, in addition to observation-based explorations, major space powers like the United States, the European Union and China have announced plans to launch execution-based missions such as sample return and base constructions. Unmanned space exploration is evolving from the one that is characterized by an individual spacecraft to the one that relies on network operations. This, however, exceeds the scope of competence of current DSNs.
To ensure that communication enhances rather than constrains unmanned space exploration, the development of space-space communication is essential for robotic cooperation. Without the challenges of long distances, space-space communication facilitates fast and flexible connections among space robots, allowing them to form intelligent cooperation systems. This alleviates the communication burden between the earth segment and the space segment. However, unlike traditional communication networks, the formed cooperation system exhibits prominent task-oriented characteristics. In most cases, space robots operate cyclically: sensors collect surrounding information and report it to the computing unit, the computing unit analyzes the situation and distributes calculated commands to actuators, and the actuators take actions. These robots synergistically form the sensing-communication-computing-control ($\mathbf{SC}^{3}$) loop, which is the basic unit for task execution. Therefore, the communication within the $\mathbf{SC}^{3}$ loop is geared toward serving a specific task rather than achieving a universal data rate.
In this paper, we explore the communication scheme for a specific cooperation system named the “mother-daughter system”. As the name suggests, the “mother-daughter system” consists of one mother spacecraft and multiple daughter probes. In Fig. 1, we illustrate the operational pattern of this system. Upon the launch of a space mission, the mother spacecraft undergoes the launch, transfer orbit, and planet approach phases before reaching the scheduled planetary orbit. Subsequently, the spacecraft releases the lander. Once the lander lands on the planetary surface, it deploys daughter probes. These daughter probes are assigned with various exploration tasks. To enable these relatively less advanced probes to conduct arduous exploration tasks, the mother spacecraft needs to provide real-time guidelines. It carries advanced sensing, communication and computing modules. In each control cycle, the mother spacecraft senses the environment, conducts analysis, and sends instructions to daughter probes to take actions. Together, they synergistically form $\mathbf{SC}^{3}$ loops, serving as the fundamental unit for task execution. Moreover, the adaptable orchestration of the “mother-daughter system” can further give rise to a large-scale exploration network.
@@
In the literature, few studies have investigated space communication from the perspective of the $\mathbf{SC}^{3}$ loop. Communication researchers mostly take the communication link as the investigated objective and address the unique challenges in DSC [5]. While improving communication performance is crucial for space exploration, separating the communication process from the $\mathbf{SC^{3}}$ loop might not be optimal. This approach overlooks the communication impact on the overall functioning of the $\mathbf{SC^{3}}$ loop. In contrast, researchers in control have studied the whole $\mathbf{SC^{3}}$ loop. Specifically, in networked control systems (NCSs), where sensors, controllers, and actuators are interconnected via a shared network, studies showed that communication imperfections can adversely affect the closed-loop control [6][8]. In the following, we review related studies of DSC and NCS.
@@ I-B1 DSC
DSC plays an important role in unmannd space exploration. Due to the long transmission distance, there exist unique challenges for DSC. Related studies have addressed different challenges from different perspectives, including communication protocols, networking, and resource optimization. For example, Ha et al. proposed a reinforcement learning-based link selection strategy to improve the streaming performance of DSC [9]. To tackle the solar scintillation effect, Xu et al. devised a dual-hop mixed communication system, which comprises a radio frequency link between the earth and the relay satellite and a free-space optical link between the satellite and a deep space probe [10]. To address the challenge of intermittent connections, the delay-tolerant network (DTN) has been developed, in which the store-and-forward strategy is used to combat the interruption. On this basis, Yang et al. proposed a hybrid bundle retransmission scheme to ensure the DTN reliability [11]. Rango et al. proposed an adaptive bundle rate scheme to tackle concurrent bundle transmissions [12]. In terms of the network structure, Wan et al. proposed a satellite relay constellation network and detailed the mathematical model, topology design and performance analysis [13]. These studies have improved the communication performance in deep space, providing a solid foundation for unmanned space exploration.
I-B2 NCS
@@
Building upon these results, numerous resource optimization schemes have been proposed for NCSs. For example, Lyu et al. devised a control-aware cooperative transmission scheme that jointly optimized channel allocation, power control, and relay cooperation [18]. Girgis et al. provided a prediction-based control scheme to minimize the age of information and transmit power [19]. Chang et al. optimized the bandwidth, transmit power, and control convergence rate to maximize the spectrum efficiency [20]. Yang et al. developed a framework to optimize a new metric named the energy-to-control efficiency [21]. Lei et al. proposed a multi-loop optimization scheme to minimize the sum LQR cost. The transmit power was balanced among different loops [22]. Esien et al. proposed a control-aware scheduling scheme that adapted transmissions to channel dynamics and system states [23]. While these works took a first step to orchestrate communication resources for the closed-loop control, they were still preliminary. Some overlooked the fact that the communication within the $\mathbf{SC^{3}}$ loop has stringent latency requirement and did not model the data transmission in the finite block length (FBL) regime, and some remain to take communication metrics as the objective.
@@ I-C Main Contributions
In this paper, we consider a space cooperation system named the “mother-daughter system”, which applies $\mathbf{SC^{3}}$ loops to execute a control-type exploration task. To improve the control performance of these $\mathbf{SC^{3}}$ loops, we propose a control-oriented communication scheme. The main contributions are listed as follows.
@@
    We propose a control-oriented communication scheme for the “mother-daughter system”. Considering the cycle-time constraint, we model the mother-daughter downlink in the FBL regime. The proposed scheme takes the LQR cost as the objective. The optimized variables are the block length and transmit power.
@@
    To solve the nonlinear integer problem, we first derive the optimal block length based on the monotonicity of the rate-cost objective function. Subsequently, we analyze the monotonicity and concavity-convexity of the achievable rate expression and the rate-cost function. Building upon these insights, we transform the power allocation problem into a tackle convex problem.
@@
    In the assure-to-be-stable region, we derive approximate closed-form solutions for the proposed scheme, as well as for two communication-oriented schemes: the max-sum rate scheme and the max-min rate scheme. By showing the equivalence between the proposed scheme and the max-min rate scheme for time-insensitive control tasks, we reveal the fairness-minded nature of the proposed scheme.
@@
    We use appropriate parameters to model the space environment in the simulation. Simulation results confirm the effectiveness of the proposed scheme in improving the control performance of the $\mathbf{SC^{3}}$ loops and validate our findings.
@@ II System Model
Figure 2: Illustration of an unmanned space cooperation system that comprises a mother spacecraft and multiple daughter probes. The spacecraft and probes synergistically form multiple $\mathbf{SC^{3}}$ loops, and each loop takes charge of a control-type subtask.
In Fig. 2, we present the model of the “mother-daughter system”. The mother spacecraft is located in the planet’s orbit and carries advanced sensing, communication and computing modules. In each control cycle, it senses system states, calculates control commands and distributes these commands to daughter probes for actions. Working in a cyclical manner, they synergically form multiple $\mathbf{SC^{3}}$ loops. The mother spacecraft also maintains a communication link with the earth, which serves for TT&C purposes and sends crucial data back to the earth. We assume that this “mother-daughter system” includes $K$ $\mathbf{SC^{3}}$ loops. Each $\mathbf{SC^{3}}$ loop operates an LQR-based controller to execute a control-type task. For the $k$ th $\mathbf{SC^{3}}$ loop at time index $i$, the following state-space equation is used to characterize the dynamics of the controlled system,
@@
where $\mathbf{x}_{k,i}\in\mathbb{R}^{m\times 1}$ is the system state and $\mathbf{u}_{k,i}\in\mathbb{R}^{m\times 1}$ is the control input. The matrices $\mathbf{A}_{k}\in\mathbb{R}^{m\times m}$ and $\mathbf{B}_{k}\in\mathbb{R}^{m\times m}$ are determined by the system dynamics, and $\mathbf{v}_{k,i}\in\mathbb{R}^{m\times 1}$ is the system uncertainty, whose covariance matrix is denoted as $\mathbf{\Sigma}_{\mathbf{v}_{k}}\in\mathbb{R}^{m\times m}$. The system is unstable $(\lambda(\mathbf{A}_{k})>1)$, and it can be stabilized using $(\mathbf{A}_{k},\mathbf{B}_{k})$. In practice, the state-space equation is non-linear. (1) is obtained by linearizing the system state around the working point. Currently, the LQR controller has been widely applied in space exploration, for tasks like spacecraft maneuvering and formation flying [24]. Although simplified, the linear control model has been proven effective in characterizing the system dynamics.
@@
Afterward, the mother spacecraft functions as the BS, sending control commands to the respective probes. We assume that both the mother spacecraft and daughter probes are equipped with a single antenna. This single-antenna setting can be extended to the multi-antenna setting, in which the control-oriented beamforming is an interesting topic. We do not delve into this case in this paper. The spacecraft uses orthogonal frequency bands to communicate with probes, allowing different $\mathbf{SC}^{3}$ loops to operate simultaneously without co-channel interference. The bandwidth is the same for $K$ loops, which is denoted as $B$. As for the channel model, the near-planet link has similar characteristics as the near-Earth link [26]. We thereby adopt the model of the near-Earth link [27] to characterize the near-planet environment:
@@
where $t_{k}^{s}$, $t_{k}^{c}$, $t_{k}^{d}$, and $t_{k}^{a}$ are the time for sensing, computing, downlink communication, and executing control actions. Since we focus on the communication process, the time for sensing, computing and executing control actions is predetermined in this paper. The communication time includes two parts,
@@
Moving forward, we use the LQR cost to measure the control performance of the $\mathbf{SC}^{3}$ loop, which is calculated by
@@
where $\mathbf{Q}_{k}$ and $\mathbf{R}_{k}$ are positive semidefinite matrices. They are used to balance the cost of the state derivation and control input. According to [15], the cycle rate needs to satisfy the following stable condition such that the controlled system can be stabilized,
@@
which establishes a bridge between the limit LQR cost and the cycle rate. This connection allows us to orchestrate communication resources for the closed-loop control, which is detailed in the next Section.

### v3

@@ I-A Background and Motivation
Communication plays a vital role in deep space exploration. The link between the spacecraft and the earth command center is the lifeline of the spacecraft. This connection provides tracking, telemetry and command (TT&C) services, which ensure the correct operation. According to the definition of the International Telecommunication Union , if the spacecraft-Earth distance is greater than $2\times 10^{6}$ km, the communication between the spacecraft and the earth belongs to deep space communication (DSC) [1]. Due to the great distance, DSC exhibits unique characteristics. The first is the large latency. Taking Mars as an example, the propagation delay from the perigee of the Mars to the Earth is 4 minutes, while from the apogee of the Mars, it is 24 minutes [2]. The second is the low data rate. As a common sense, the signal power decays at least quadratically with distance. To ensure the same strength of the received signals, the transmit power of the spacecraft from Mars should be $9\times 10^{16}$ times stronger than that of a smart phone located 1 km away from the receiver [2]. However, constrained by size and weight limitations, the on-board base station (BS) carried by the spacecraft is less powerful than ground-based facilities. It can only transmit signals at very low power. Ground receivers must be equipped with large aperture antennas to capture the weak signals from space. Even so, the data rate between the spacecraft and the earth command center remains quite low, e.g., several megabits per second in NASA’s Deep Space Network (DSN) [3].
@@
With the growing curiosity on outer space, current space activities touch the fundamental problems of DSC. The sheer number of space missions is one factor. According to the United Nations, the number of space robots has grown exponentially in the last decade [4]. With advanced sensing techniques, new spacecrafts need a higher data rate to send high-resolution data back, which dramatically increases the burden of the DSN. Moreover, in addition to observation-based explorations, major space powers like the United States, the European Union and China have announced plans to launch execution-based missions such as sample return and base constructions. Unmanned space exploration is evolving from the one that is characterized by an individual spacecraft to the one that relies on network operations. This, however, exceeds the competence of current DSNs.
Unlike space-ground communication that is suffered from long distances, the space-space communication enables fast and flexible connections among space robots, fostering robotic cooperation. With the connection of the communication links, space robots operate cyclically: sensors collect surrounding information and report it to the computing unit, the computing unit analyzes the situation and distributes calculated commands to actuators, and the actuators take actions. These robots synergistically form sensing-communication-computing-control ($\mathbf{SC}^{3}$) loops. However, communication often becomes the critical bottleneck in the $\mathbf{SC}^{3}$ loop. Firstly, to deal with the uncertain environment, the $\mathbf{SC}^{3}$ loops run in small cycle time. This constrains the time for data transmissions. When the spacecrafts use the short packet to transmit information, the achievable rate suffers significant degradation compared with the traditional Shannon capacity. In addition, the BSs carried by the spacecraft are less powerful compared with ground facilities. This further limits the communication data rate. To ensure that communication enhances rather than constrains unmanned space exploration, addressing communication challenges emerges as a pivotal factor in enhancing the $\mathbf{SC}^{3}$ loop.
In this paper, we explore the communication scheme for a cooperation system named the “mother-daughter system”. As the name implies, this system comprises a central robot serving as the “mother”, alongside multiple less advanced robots functioning as the “daughters”. To facilitate daughter robots’ tasks, the mother robot diligently oversees daughter robots and provides them with sensing and computing assistance and operational guidelines. In fact, the “mother-daughter system” is not a new concept. We are able to find many prototypes of such systems in current space missions. In Mars 2020, the Perseverance rover (“mother”) and the Ingenuity helicopter (“daughter”) is a typical example. Throughout Ingenuity’s missions, Perseverance monitored its flight and relayed data between the Ingenuity and the Earth control center. Supported by the Perseverance, the Ingenuity navigated numerous unforeseen challenges beyond its initial design, including choosing landing sites in treacherous terrain, dealing with a dead sensor, and cleaning itself after dust storms [5]. Another example is the Lunar gateway, which is a vital component of the NASA-led Artemis missions planned to be launched in 2025. This space station will be a multi-purpose outpost orbiting the Moon, acting as a “mother” space station to support “daughter” landers, rovers, and human missions on the Moon’s surface. Such configurations, involving a central spacecraft (“mother”), coordinating with multiple surface or subsurface probes (“daughters”), are increasingly recognized as a basic coordination pattern for the unmanned space exploration. In Fig. 1, we illustrate the envisaged “mother-daughter system” in future unmanned space missions. Upon the launch of a space mission, the mother spacecraft progresses through launch, transfer orbit, and planet approach phases to reach its designated orbit. It then releases the lander, which descends to the planetary surface and releases daughter probes. To enable these less advanced probes to conduct exploration tasks in varying environments, the mother spacecraft provides real-time guidelines to them. Together, they form $\mathbf{SC}^{3}$ loops, serving as the fundamental unit for task execution. In addition, the adaptable orchestration of the “mother-daughter system” can further give rise to a large-scale exploration network.
In the literature, few studies have investigated space communication from the perspective of the $\mathbf{SC}^{3}$ loop. Communication researchers mostly took the communication link as the investigated objective and addressed the unique challenges in DSC [6]. While improving communication performance is crucial for space exploration, separating the communication process from the $\mathbf{SC^{3}}$ loop might not be optimal. This approach overlooks the communication impact on the overall functioning of the $\mathbf{SC^{3}}$ loop. In contrast, researchers in control have studied the whole $\mathbf{SC^{3}}$ loop. Specifically, in the wireless control system (WCS) [7, 8, 9], which is a type of distributed internet of things (IoT) system where sensors, controllers, and actuators are interconnected via a wireless network, communication researchers began to take the $\mathbf{SC^{3}}$ loop as the research objective and proposed the control-aware scheme. In the following, we review related studies of DSC and WCS.
@@ I-B1 DSC
DSC plays an important role in unmannd space exploration. Due to the long transmission distance, there exist unique challenges for DSC. Related studies have addressed different challenges from different perspectives, including communication protocols, networking, and resource optimization. For example, Ha et al. proposed a reinforcement learning-based link selection strategy to improve the streaming performance of DSC [10]. To tackle the solar scintillation effect, Xu et al. devised a dual-hop mixed communication system, which comprises a radio frequency link between the Earth and the relay satellite and a free-space optical link between the satellite and a deep space probe [11]. To address the challenge of intermittent connections, the delay-tolerant network (DTN) has been developed, in which the store-and-forward strategy is used to combat the interruption. On this basis, Yang et al. proposed a hybrid bundle retransmission scheme to ensure the DTN reliability [12]. Rango et al. proposed an adaptive bundle rate scheme to tackle concurrent bundle transmissions [13]. In terms of the network structure, Wan et al. proposed a satellite relay constellation network and detailed the mathematical model, topology design and performance analysis [14].
In addition to these straightforward communication schemes, some cutting-edge technologies can further enhance the deep space network. For instance, the non-orthogonal multiple access (NOMA) technique can be introduced to allow multiple space robots to share the same spectrum resources. Its application in DSC could enable more efficient usage of spectrum resources. Moreover, equipping space robots with advanced and compact hardware, such as reconfigurable pixel antennas and integrated communication and sensing platforms, could enhance their operational capabilities. Furthermore, the age of information (AoI) provides a new perspective to measure the freshness of data. Recently, Yuan et al. proposed an AoI-guided user scheduling scheme for the communication between the Lunar far surface and the Earth. The proposed scheme was proven effective to ensure the timely delivery of sensing data [19]. Thinking more broadly, the unmanned aerial vehicle (UAV) can be uploaded to the explored planet and establish the satellite-UAV-ground network to felicitate large-scale unmanned space exploration [15][16]. These techniques provide valuable insights for enhancing the deep space network and deserves more research attentions.
I-B2 WCS
@@
Building upon these results, numerous control-aware schemes have been proposed for WCSs. Weiner et al. proposed a control-aware wireless system and designed the transmission schemes in physical and media access control layers to guarantee the latency of the command transmission [22]. Gatsis et al. proposed a sensor transmission policy to minimize the LQR cost while constraining of the long-term communication resource consumption [23]. Lyu et al. devised a control-aware cooperative transmission scheme that jointly optimized channel allocation, power control, and relay cooperation [24]. Tong et al. considered a predictive-based control system, which transmits a set of predicted commands in a single transmission. The authors optimized the predictive command length to minimize the communication consumption [25]. Girgis et al. provided a prediction-based control scheme to minimize the AoI and transmit power [26]. Chang et al. optimized the bandwidth, transmit power, and control convergence rate to maximize the spectrum efficiency [27]. Yang et al. developed a framework to optimize a new metric named the energy-to-control efficiency [28]. Lei et al. investigated a satellite-UAV system and proposed a multi-loop optimization scheme to minimize the sum LQR cost. The transmit power was balanced among different loops [29]. Esien et al. proposed a control-aware scheduling scheme that adapted transmissions to channel dynamics and system states [30]. Recently, Ali et al. incorporated the deep learning approach to optimize the sampling period, communication block length, and packet error probability to minimize the energy consumption [31]. While these works took the first step to investigate the control-aware communication design, challenges remain in the considered space setting. Firstly, the space environment is totally uncertain. The $\mathbf{SC}^{3}$ loop needs to run in small cycle time to adapt to the varying environments. This restricts the time window of data transmissions. When using short packets in communication, the achievable rate in the finite block length (FBL) regime is largely reduced compared with the Shannon capacity. Moreover, given the payload constraint, space robots are equipped with lightweight transceivers, whose communication capabilities is far away from ground-based facilities. The time sensitivity of the $\mathbf{SC}^{3}$ loop and limited communication resources of the space robots call for the accurate modeling and efficient usage of communication in space missions. Current studies either failed to prioritize control metrics as their objectives or did not model the data transmissions in the FBL regime, making them lose efficiency in the space environment. Tab. I presents a qualitative comparison of this work and related work in WCSs.
TABLE I: Quantitative Comparisons between the proposed scheme and the related works in WCSs
Work System Setting Optimization Objective Short-packet Transmission Loop Number
Weiner et al. [22] Real-time WCS Communication: latency Yes Multiple loops
Gatsis et al. [23] General WCS Control: LQR cost No Single loop
Lyu et al. [24] Multi-hoop WCS Hybrid: transmission ability and the upper bounds of the LQR cost No Multiple loops
Tong et al. [25] Pridiction-based WCS Communication: energy consumption No Single loop
Girgis et al. [26] General WCS Hybrid: a weighted summation of the AoI and the transmission power No Multiple loops
Chang et al. [27] Real-time WCS Communication: energy consumption Yes Single loop
Yang et al. [28] Real-time WCS Hybrid: energy-to-control efficiency Yes Multiple loops
Lei et al. [29] Satellite-UAV system Control: limit LQR cost No Multiple loops
Esien et al. [30] Real-time WCS Communication: transmission time Not covered Multiple loops
Ali et al. [31] Real-time WCS Communication: power consumption Yes Multiple loops
This work Mother-daughter system in space Contol: limit LQR cost Yes Multiple loops
Recently, there are some works that considered the joint communication and sensing/computing model in the closed-loop control. Han et al. introduced the reconfigurable intelligence surface to cover the blind areas of the WCS. The authors introduced the rate-distortion function to model the sensing phase, which imposes the uplink to transmit a certain quantity of data to meet the distortion requirement. The proposed scheme remained a communication scheme that jointly optimized the transmission power, transmission time, beamforming and the reflecting coefficients of RIS [32]. Li et al. proposed a joint communication and computing optimization scheme, which took the communication power and computing frequency as the optimization variable and minimized the closed-loop latency [33]. These works made initial attempts to combine sensing, communication, and computing within the $\mathbf{SC^{3}}$ loop modeling. However, a foundational model that can uniformly characterize the relationships among sensing, communication, computation, and control is still unknown. In the considered space setting, we identify communication as the critical bottleneck affecting the overall performance of the $\mathbf{SC^{3}}$ loop. Given the limited communication capabilities of space robots, maximizing the communication efficiency for the task becomes crucial in enhancing the $\mathbf{SC^{3}}$ loop. Therefore, our paper focuses on the communication resource optimization for the closed-loop control. We take the joint sensing, communication, and computing design as an important direction for future research.
@@ I-C Main Contributions
In this paper, we investigate the communication scheme for the $\mathbf{SC^{3}}$ loop within the “mother-daughter system”. Given the task-oriented trait of the $\mathbf{SC^{3}}$ loop, we propose a control-oriented communication scheme, whose distinctive feature is the integration of control objective into the communication design. The main contributions are listed as follows.
@@
    We propose a control-oriented communication scheme for the $\mathbf{SC^{3}}$ loop within the “mother-daughter system”. Considering the cycle-time constraint, we model the mother-daughter downlink in the FBL regime. On this basis, we optimize the downlink block lengths and transmit power to minimize the sum LQR cost of the “mother-daughter system”.
@@
    To solve the nonlinear integer problem, we first derive the optimal block lengths based on the monotonicity of the rate-cost objective function. Subsequently, we rigorously prove the monotonicity and concavity-convexity of the achievable rate expression and the rate-cost function. Building upon these insights, we transform the power allocation problem into a tackle convex problem.
@@
    Leveraging the tight approximations of the cycle rate and rate-cost functions, we derive the approximate closed-form solution of the transmit power. We compare this closed-form solution with that of the max-sum rate scheme and the max-min rate scheme. Through comparative analysis, we differentiate the proposed scheme with communication-oriented schemes and find its fairness-mined nature.
@@
    We use appropriate parameters to stimulate the space environment and test the proposed scheme. Simulation results confirm the effectiveness of the proposed scheme in improving the control performance of the $\mathbf{SC^{3}}$ loops and validate our findings.
@@ II System Model
Figure 2: Illustration of an unmanned space cooperation system that comprises a mother spacecraft and multiple daughter probes. The spacecraft and probes synergistically form $\mathbf{SC^{3}}$ loops, and each $\mathbf{SC^{3}}$ loop takes charge of a control-type subtask.
In Fig. 2, we present the model of the “mother-daughter system”. The mother spacecraft is located in the planet’s orbit and carries advanced sensing, communication, and computing modules. In each control cycle, it senses system states, calculates control commands and distributes these commands to daughter probes for actions. Working in a cyclical manner, they synergically form multiple $\mathbf{SC^{3}}$ loops. The mother spacecraft also maintains a communication link with the Earth, which serves for TT&C purposes and sends crucial data back to the Earth. We assume that this “mother-daughter system” includes $K$ $\mathbf{SC^{3}}$ loops. Each $\mathbf{SC^{3}}$ loop operates an LQR-based controller to execute a control-type task. For the $k$ th $\mathbf{SC^{3}}$ loop at time index $i$, we use the following linear time-invariant equation to model the dynamics of the controlled system,
@@
where $\mathbf{x}_{k,i}\in\mathbb{R}^{m\times 1}$ represents the system state, $\mathbf{u}_{k,i}\in\mathbb{R}^{m\times 1}$ represents the control actions, and $\mathbf{v}_{k,i}\in\mathbb{R}^{m\times 1}$ represents the system uncertainty, whose covariance matrix is denoted as $\mathbf{\Sigma}_{\mathbf{v}_{k}}\in\mathbb{R}^{m\times m}$. The matrix $\mathbf{A}_{k}\in\mathbb{R}^{m\times m}$ is the system dynamics matrix, which quantifies the intrinsic behavior of the states in the absence of control inputs. The matrix $\mathbf{B}_{k}\in\mathbb{R}^{m\times m}$ signifies the control input, delineating how control actions ($\mathbf{u}_{k,i}$) influence the system state transition from $\mathbf{x}_{k,i}$ to $\mathbf{x}_{k,i+1}$. Taking the temperature control as an example, the system state is thus the temperature. $\mathbf{A}_{k}$ models how the current temperature evolves into the next value without outside intervention and $\mathbf{B}_{k}$ models the impact of the control action on the temperature changes. Back to our setting, we assume the system is unstable $(\lambda(\mathbf{A}_{k})>1)$, and it can be stabilized using $(\mathbf{A}_{k},\mathbf{B}_{k})$. In practice, the state-space equation is non-linear. (1) is obtained by linearizing the system state around the working point. Currently, the LQR controller has been widely applied in space exploration, for tasks like spacecraft maneuvering and formation flying [34]. Although simplified, the linear control model is effective in characterizing the system dynamics.
@@
Afterward, the mother spacecraft functions as the BS, sending control commands to the respective probes. We assume that both the mother spacecraft and daughter probes are equipped with a single antenna. This single-antenna setting can be extended to the multi-antenna setting. Notably, since the channel in space environment is less affected by small-scale fading, the multi-antenna configuration is utilized to implement beamforming. By concentrating the signal energy towards the intended receiver, beamforming significantly increases the signal-to-noise ratio (SNR). This improvement in the SNR is crucial for our “mother-daughter system”, where the signal strength can be severely attenuated due to the vast distance of the mother-daughter downlink. In addition to enhancing the SNR, directional communication helps minimize inter-link and inter-system interference. This ensures the performance of the whole space network, accommodating the demands of large-scale exploration in the future. We take the control-oriented beamforming as an interesting topic for further research and we do not delve into this topic in this paper. The spacecraft uses orthogonal frequency bands to communicate with probes, allowing different $\mathbf{SC}^{3}$ loops to operate simultaneously without co-channel interference. The bandwidth is the same for $K$ loops, which is denoted as $B$. As for the channel model, the near-planet link has similar characteristics as the near-Earth link [36]. We thereby adopt the model of the near-Earth link [37] to characterize the near-planet environment:
@@
where $t_{k}^{s}$, $t_{k}^{c}$, $t_{k}^{d}$, and $t_{k}^{a}$ are the time for sensing, computing, downlink communication, and executing control actions. Since we focus on the communication process, the time for sensing, computing, and executing control actions is predetermined in this paper. The communication time includes two parts,
@@
Moving forward, we use the LQR cost to measure the control performance of the $\mathbf{SC}^{3}$ loops. It is a weighted quadratic summation of the state derivations and the control inputs,
@@
where $\mathbf{Q}_{k}$ and $\mathbf{R}_{k}$ are weight matrices. $\mathbf{Q}_{k}$ quantifies the importance of maintaining state variables within desired values and $\mathbf{R}_{k}$ quantifies the relative cost or undesirability of utilizing control inputs to influence the system state. As a result, the larger $\mathbf{Q}_{k}$ leads to more powerful control actions, while the larger $\mathbf{R}_{k}$ leads to more cost-friendly control actions. In practical implementations, engineers adjust $\mathbf{Q}_{k}$ and $\mathbf{R}_{k}$ to produce a controller consistent with designed goals. In the context of our “mother-daughter system”, the LQR cost represents a theoretical measurement of the control efficiency, with a lower LQR cost denoting more effective and efficient control of the $\mathbf{SC}^{3}$ loop. When it is applied in practical applications, the LQR cost is mapped to a specific physical meaning. For instance, in regulating extreme temperatures, which is the basis for many space experiments, e.g., space breeding, the LQR cost quantifies temperature deviations from the desired range and the cost of control inputs. A lower LQR cost represents the temperature is more effectively maintained in the desired interval. Thus, compared with general communication metrics, e.g., rate, the LQR cost has more specific physical meanings and can directly reflect task performance. This is why we take the LQR cost as the objective in this paper.
According to [18], the cycle rate needs to satisfy the following stable condition such that the controlled system can be stabilized,
@@
which establishes a bridge between the limit LQR cost and the cycle rate.